OpenStudy (anonymous):

write the standard form of the equation of the line that is parallel to the graph of the given equation and that passes through the point with the given coordinates

OpenStudy (anonymous):

4. 2x+2-5=0; (0,4) 5. 3x-y+3=0; (-1,-2)

OpenStudy (triciaal):

parallel lines have the same slope

OpenStudy (triciaal):

given 2 points on the line (x,y) and (x1, y1) and the slope of the line, the equation of the line is (y-y1) = m(x-x1) where m is the slope

OpenStudy (triciaal):

standard form of the equation of the line Ax + by = C